OpenStudy (anonymous):

Suppose the probability of an event occurring is P(A), and the probability of the event not occurring is P(A'). If P(A) = m and P(A') = n, which of the following equations must be true? A. B. C. D.

OpenStudy (campbell_st):

well as there are no choices shown...I'll guess it will be something like m + n = 1 or m = 1 -n or n = 1 - m